Rewrite the equation 2x – 3y = -6 as a funcion of x.
Answer by fractalier(6550)

y can be expressed as a function of x in this, and in most, examples at this level...thus we merely solve for y...
2x – 3y = -6
-3y = -2x - 6
y = (2/3)x + 2 or
f(x) = (2/3)x + 2